Grade 12Physical Chemistry

The ratio of wavelength of de brogeli wavelength of the electron in the first and third orbit of hydrogen atom

Abhilasha
9 Years agoGrade 12

1 Answer

Vikas TU
9 Years ago
From nohr's angular momenta conservation we know that,
mvr = nh/2pi
Then For 1st orbit,
2*pi*r = 1*k1
For 2nd orbit,
2*pi*r = 2*k2
Ratio would be=> k1/k2 => 2/1 => 2:1